A man pushes a box of 40kg up incline of 15⁰, if the man applies a horizontal force 200N and the box moves up the plane a distance of 20m at a constant velocity and the coefficient of friction is 0.10, find
1) work done by the man on the box
2)work done against friction
1
Expert's answer
2021-08-24T06:53:15-0400

1) The work done by the man


"W=F\cdot s \cdot\cos\alpha=200\cdot20\cdot\cos 15°\approx3864\ (J)"


2) The work done against friction


"W=F_f\cdot s =(mg\cdot \cos15°+200\cdot\sin 15°)\cdot k\cdot s="


"=(40\cdot9.81\cdot \cos15°+200\cdot\sin 15°)\cdot 0.1\cdot 20\approx861\ (J)"